50 recent changes in AdaptiveOptics Web retrieved at 21:09 (GMT)

20240228_LBTIDX_Day
Details. Power on DX. OK. OCAM OK. Closeloop test. OK. Fernando has to go back home. Rejoin in 2 Hours. Running the loop with the soul version for zero slopes . W...
20240219_SXDX_Day
Staff: lbto: Juan Carlos Observer: NA Summary. * Continuing testing the zero slope firmware upgrade for soul lucisx and dx. The test has not shown any zero sl...
20240218_LBTIDX_shark
SX SHARKNIR log can be found here: https://docs.google.com/document/d/13EG_6UKz7UAde9J9TBIkjXcI2QUnlyQp5mDD 52U8wA/edit?usp=sharing Forecast. Details. Powe...
20240218_SX_Night_UT
SX LUCI AO 1:46 * Closing the loop. Seeing isn't great, but there is a large delta T between the primaries and the outside air. 2:16 * Adsec not sending GOP...
20240218_DX_Night
Staff: Lbto: Josh, Justing, Greg, Juan Carlos Observer. INAF. Remote Italy Forecast. Details. Power on 05:30PM Ocam OK. 05:53PM Justin just communicate that it ...
20240216_lbtidx_Night
Staff: LBTO: Josh, Justin, Greg, Juan Carlos Observer: Sharkvis (Fernando,Roberto). Remote Italy. SharkNir (). Remote Italy. INAF. remote Italy Forecast. Detail...
20240217_LBTISX_Night_UT
SHARK NIR 1:14 * Open * IE 1.0 CA 50.0 * Current position is X=66.487, Y=17.819Current position saved into fov.conf as "shark"File fov.txt has been updat...
AsgardGui
Asgard * Asgard as of 07nov2021 * Asgard as of 23nov2023 * material.io
20240216_SXDX_Day
Details. Version deployed Currently running: test.20240125.zero slope firmware (soul) Power on SX. ocam with bad. TN:20240216_195420 power cycle bcu2k preset DX F...
20240215_LBTDX_Night
Staff. lbto: Josh, Justin, Juan CArlos, greg Observer: Fernando (shark vis). RE mote Italy. Shark Nir (luca). remote Italy. INAF. Elena. Remote Italy Forecast. ...
20240216_LBTISX_Night_UT
SHARK NIR 1:30 * CQ Tau * Bin 1 833 Hz Mag 10.7 * Everything is running smoothly. * sxwunit@lbti sxwfs source $ fov_select.py s sharkConfiguration s...
20240214_LBTIDX_Night
Staff: LBTO: Josh,Justin, Juan Carlos, Greg (lbtiSX) Observer: Shark_vis team. Remote Italy. Shark nir team . remote Italy Forecast. Details. Same target.LKCal...
20240214_LBTISX_Night
SHARK NIR IE 3.0 CA 54.0 1:59 * LkCA15 * Bin 2. 823 Hz * Going to take a LOOONG dataset * 20240215_020033 * 20240215_020249 * 20240215_020856 ...
20240213_LBTIDX_Night
Forecast. Details. Night log: https://docs.google.com/document/d/1c2k5IBfSww_hqc9DgcVRurcwWENkrQZjZ27Kdi5KYDQ/edit email from Simone (sharkvis): " as you know SV...
20240209_SX_Day
Details. Power on. WfsArbitrator_35623 INF 66 2024 02 09 15:21:52.161635 MAIN Cannot turn on FAN_47. Reason: Timeout waiting for powerboard.L...
20240208_SX_Day
Details. 09:00 Power on Ocam OK 10:00 Power bump. adsec in failure. FLAO message. 2024 02 08 16:57:20.966969 Error2024 02 08 16:57:30.684584 13006 AOExcepti...
20240208_DX_Evening
Purpose, to determine the focus offset between LBTI and SNIR. Tested dichroic thoroughput difference on sky on UT 2023 Dec 28 https://docs.google.com/spreadsheets...
20240207_SX_Day
Staff: LBTO: Juan Carlos Observer: Summary. Close loop test using the zero slope firmware. Acquire performances images with Luci and test the ncpa measurements. ...
20240206_SX_Day
Details. Power on from AOS. WfsArbitrator_36993 INF 75 2024 02 06 15:36:22.316059 MAIN Missing network connection with ts8 dx. Check network ...
Log_pointing
Pointing and Collimation For SOUL LBTI WFS Focal Station LBTI DATEELEVATIONAUTHORIZATIONINSTRUMENT MODESX IE/CADX IE/CATEMPNOTE 2023 09 29 N/A Bino LBTI LMIRC...
20240202_SX_Day
Staff: LBTO: Juan CArlos Observer. Summary. Test of the new firmware for the zero slopes events. No events reported during all the time that the loop was closed....
20240202_DX_Day
Delta Automation Tests 17:19 * Source Found, Loop closed, everything optimized OK. * Hexapod: 4.1 1.9 0.705 177.8 239.4 18:14 * This works, not sure h...
20240131_SX_Day
Staff: lbt0: Brandon, Juan CArlos Observer: Fabio (remote Italy) Details. link of notes: https://docs.google.com/spreadsheets/d/1WTxT8pAMRPsS7YU_kSXqWBFzt3kcclbO...
202401126_SX_day
Staff: LBTO: Brandon, Juan CArlos Observer Details. test plan: https://docs.google.com/spreadsheets/d/1WTxT8pAMRPsS7YU_kSXqWBFzt3kcclbO5mlcD8T0xqw/edit#gid=2138...
20240125_X_DAy
Staff: LBTO: Brandon, Juan Carlos Observer: Fabio (Arcetri rremote) Summary. Test the zero slope firmaware update from mario and checkout that the correct softwa...
Remove_bad_actuator_new
To remove a bad actuator requires an identificaiton of the bad actuator, modify the bad_actuators.list and execute and IDL script. * Check in the AdSec Mirror ...
Procedures
Remove Bad or jump actuator. Remove Bad or jump actuator (old). Clear daytime activity. Flash BCU47 firmware. Deployment Soul version environments. Calculate...
20240123_SXDXLBTI_Day
Staff: LBTO: Juan Carlos Observer. Details. Chamber temp 6.56 Power on soul lucisx ocam ok Change to zero slope version. not events recorded. Startup LBTI WFSs. ...
NotLive
OCAM Bad Frames * JP 25sep2023 email * JP 22jan2024 meeting minutes
20240119_DX_Day
Details. 09:00am Power on Ocam ok Delayed the work becaue the ICS was rechargng and the Adsec was not ready. 10:20PM Adsec ready. Task#1. Close loop Automatic mo...
20240119SXDX_NIGHT_UT
LBTB 1:00 * Open and ready to go. 1:26 * Running ESM Bino Mag 13.1 SX 12.6 DX 2:50 * Seeing has really exploded (it's 1.5", but I saw 2.5"), but ESM is chu...
20240117_SXDX_Night
Details. Observing logs. https://docs.google.com/document/d/10r8MWnW4TbOFVyYeDQqbUwAcS5cbaMhP U_SvOQ5P_I/edit?usp=sharing_eil_m ts=65a86995 05:00PM Power on WFSs ...
20240117_DX_Day
Details. There are some work on MODS troubleshooting. Lights on in the chamber. Working with lower lights on. Power on ocam ok 11:30 change version: Currently run...
20240116_SXDX_Night
Staff: LBTO: David,Andrew, Juan Carlos Observer: Don,Michael (OSU) (remote) forecast. Details. 04:30pm Power on DX/SX OK. Ocam SX/DX OK 05:21PM ocamsxdx cold. ...
20240116_SXDX_Day
Staff: LBTO: Juan Carlos and Brandon. Observer: Fabio (remote Italy) Summary. Details. Starting 09:45 Brandon compile and install a new environment version. Cur...
20230112_DX_Day
Staff: Lbto: Juan carlos Observer: Fernando, Simone (Shark vs team). Steward remote control room Summary Details. power on 10:00 ocam ok test close loop 1300 ga...
20240110_DX_Night
SHARK VIS 1:23 * Had a failure with the BCU2K that was causing the loop to go not live. * Resetting the BCU2K was the fix. * This happened on an attempte...
20231230_SX_Night
SHARK NIR * Opened and running. * Getting cloudier * IE 3.0, CA 63.5 03:29 * Going to take some seeing test data. * 20231230_033048 : DIMM unavail...
20231231_SX_Night
SHARK NIR 0:14 * Waiting for sunset and some clouds 2:38 * Had to cycle the AdSec to fix a fastlink error. * Going to first science target, once we sort ...
AcceptanceReviews
* SOUL SX LBTI Acceptance Review * SOUL DX LBTI Acceptance Review * SOUL DX LUCI Acceptance Review
20240110_DX_Day
Details. chamber temp 3.15 10:00 power on WFs. Ocam ok. CCD47 OK. dichroic flux test. PSF on pyramid. 11.9 400 700 flux. 12.3 window flux. 12.9 running re write ...
20240108_DX_Day
Staff: LBTO: Juan CArlos, Sam, doug, John Observer: SHARK VIS Team (telescope control room) Details. Authorize shark. Chamber temperature 5.1C check filters fl...
20230105_SXDX_Day
Details. Power on SXDX. Ocam SX OK. AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0 0 ,soy:0,rox:0 0 ,roy:0,mag:10,cindex:0,,elev:nan,rotang:nan,grvan...
20230104_SXDX_Day
Summary. Close loop withg disturbance for both Soul LuciWFS. Power bump and tume to recover limited the time of close loop. Close loop without lucis and disturban...
20240101_SXLBTI_Night
Staff: Lbto: David, Andrew,Juan Carlos Obsever:Justin (Lbti), Shark nir team on site. Forecast. Details. HAPPY NEW YEAR....ALL THE BEST. Power on SX. Ocam OK ...
20231231_SXLBTI_Night
Staff: LBTO: David, Andrew, Jared, Juan Carlos. Observer: Remote INAF (shark nir team) Forecast. Details. Power on WFS. Sun Dec 31 22:16:49.906 2023 aos.info ...
20231219_DXSX_Night
Staff: LBTO: Steve, Justin, Juan CArlos Observer: Queue Forecast. Details. 04:10 Power on SX/DX. SX. Ocam OK Dx Ocam ok cloud cover. Stay closed. WfsArbitrator ...
20231221_SXDX_Night
Staff: lbto: Steve,Alex,Juan CArlos Observer; LBTB Remote. Forecast. Details. 03:45 Power on WFSs DX. Omca ok SX. Ocam ok 05:42 Not open due to high humidity. S...
20231220_SXDX_Night
Forecast. Details. 04:20 Power on. DX WfsArbitrator_33514 INF 56 2023 12 20 23:20:01.030316 MAIN Turning off RESET_BCU_47WfsArbitrator_33514...
LoopOpen_Reclose
Description: The loop is closed and an event triggers an open loop. The star still in the pupils and the operator tries to use the FLAO gui to reclose. The reclos...
Number of topics: 50
< Previous Page 2 of 34 Next >

See also: rss-small RSS feed, recent changes with 50, 100, 200, 500, 1000 topics, all changes
Topic revision: r1 - 26 Feb 2018, ProjectContributor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback